OpenStudy (akgorgeous):

ALGEBRA QUESTION- What is the equation of the line that passes through the points (−2, 1) and (1, 10)?

OpenStudy (akgorgeous):

3x − y = −7 x + 3y = −5 3x − y = −5 3x − y = 5

OpenStudy (akgorgeous):

@3mar

OpenStudy (3mar):

Well, I am here. Are you familiar with that?

OpenStudy (akgorgeous):

yes

OpenStudy (3mar):

so can you plug in your two points into that formula and write the equation of this line? I am with you if you faced any difficulties!

OpenStudy (3mar):

@akgorgeous Do you still need help?

OpenStudy (akgorgeous):

no thankyou

OpenStudy (3mar):

Don't mention it! As you wish... I hope you got the idea!